A circle is inscribed in a triangle with sides of lengths 3 cm, 4 cm, and 5 cm. The radius of this circle, in cm, is:
(A) 5/7 (B) 1 (C) 7/5 (D) 2 (E) none of these

Answer by Alan3354(21605)

r = sqrt [(s - a)(s - b)(s - c)/s] where s = (a + b + c)/2
-----------
s = (3+4+5)/2 = 6
r = sqrt(3*2*1/6)
r = 1 cm
------------
This is very similar to Heron's Law for the area of a triangle, the difference being that "s" inside the radical is a divisor rather than a multiplier.
It works for all triangles, not just right triangles.